OpenStudy (veeveeniko):

What is the future value to the nearest dollar if you invest $25,000 today and earn 9% compounded monthly for 20 years?

OpenStudy (misty1212):

HI!!

OpenStudy (veeveeniko):

hello

OpenStudy (misty1212):

try \[25,000\times \left(1+\frac{.09}{12}\right)^{240}\] with a calculator

OpenStudy (veeveeniko):

Alright

OpenStudy (misty1212):

or just use this http://www.wolframalpha.com/input/?i=25000%281%2B.09%2F12%29%5E%28240%29

OpenStudy (veeveeniko):

I got 161,883.92

OpenStudy (veeveeniko):

wait i did it wrong

OpenStudy (veeveeniko):

I got 150,228.79